OpenStudy (anonymous):

Glen is looking into buying a piece of land. He estimates that the land measures 100 yd by 850 yd. The land is actually 140 yd by 940 yd. What is the percent error in the estimated area of the land?

OpenStudy (anonymous):

calculate the 2 areas then use the % error equations which is | measured - actual| /actual *100

OpenStudy (anonymous):

@completeidiot is right, do you understand his working?

OpenStudy (anonymous):

so add 140yd + 940yd together?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

nvm i think i got it..

OpenStudy (anonymous):

kk, tell me what you get in the end.

OpenStudy (anonymous):

100+850=950 140+940=1080 1080-950=130 im stuck after this.. is this right ?

OpenStudy (ganpat):

so (130/1080 ) * 100

OpenStudy (ganpat):

@JayDS : Should`nt it be area calculation ??

OpenStudy (anonymous):

yes, u are right, @buddhababy002 did not take note of it.

OpenStudy (anonymous):

@buddhababy002 is should rather be 100x850=85000 140x940=131600 131600-85000=46600

OpenStudy (anonymous):

now can you continue on from there?

OpenStudy (anonymous):

okay.. let me try that again and see if i can do it right..

OpenStudy (anonymous):

and in case you didn't notice @buddhababy002, we times because the area is figured out by (lxw) in this case (100x850) for the measured measurements

OpenStudy (anonymous):

kk, take your time.

OpenStudy (anonymous):

so after 131600-85000=46600 do i divide 46600 by 131600? Cause i did that and got 2.824??

OpenStudy (anonymous):

in the equation up where completeidiot wrote: calculate the 2 areas then use the % error equations which is | measured - actual| /actual *100 actual/actual *100 wouldn't that be 85000/85000*100? I'm lost....

OpenStudy (anonymous):

nono.

OpenStudy (anonymous):

\[(\frac{measured-actual}{actua}l ) \times 100\]

OpenStudy (anonymous):

actual* for the denominator, and sorry I'm having trouble with my internet.

OpenStudy (anonymous):

but the top is measured-actual bottom is actual then times the fraction after you calculate it by 100 to get the final answer as the percentage error

OpenStudy (anonymous):

I also have to go, so @Ganpat will help you out if you still need it.

OpenStudy (ganpat):

yes sir, Roger that !!!

OpenStudy (anonymous):

so would it be like this 131600-85000/131600*100?

OpenStudy (ganpat):

actual area = 140 * 940 = 131600 sq ft Measured error = 100 * 850 = 85000 sq ft.. error = 131600 - 85000 = 46600 sq ft.. so % error = (error/ actual) * 100 = (46600/ 131600) * 100 = 0.3541 * 100 35.41 % is error in measuring..

OpenStudy (anonymous):

yeah, i just finally figured out what i was doing wrong.. lols but thank you i really appreciate you explaining it more in details.

OpenStudy (ganpat):

ha ha, i guess it was @JayD who was in much detail.. i was in brief.. You should thank him..